[INSTRM-342] CCD temperature readings are nuts, Created: 30/Apr/18  Updated: 13/Jul/18  Resolved: 13/Jul/18

Status: Won't Fix
Project: Instrument control development
Component/s: ics_xcu_fee
Affects Version/s: None
Fix Version/s: None

Type: Task Priority: Normal
Reporter: cloomis Assignee: shope
Resolution: Won't Fix Votes: 0
Labels: None
Remaining Estimate: Not Specified
Time Spent: Not Specified
Original Estimate: Not Specified


 Description   

The readings from the CCDs went bad during the cool down. For both, the warmest good reading was 144.38. Below that all readings have been crazy:

2018-04-30 07:35:12.180Z cmds             20 Actor.py:491 new cmd: temps status
2018-04-30 07:35:12.181Z root             10 feeControl.py:798 sending command :'~rt,FEE\n':
2018-04-30 07:35:12.184Z root             10 feeControl.py:856 received :~rt,FEE:
2018-04-30 07:35:12.242Z root             10 feeControl.py:856 received :291.50:
2018-04-30 07:35:12.242Z root             10 feeControl.py:798 sending command :'~rt,PA\n':
2018-04-30 07:35:12.246Z root             10 feeControl.py:856 received :~rt,PA:
2018-04-30 07:35:12.447Z root             10 feeControl.py:856 received :270.63:
2018-04-30 07:35:12.447Z root             10 feeControl.py:798 sending command :'~rt,ccd0\n':
2018-04-30 07:35:12.451Z root             10 feeControl.py:856 received :~rt,ccd0:
2018-04-30 07:35:13.119Z root             10 feeControl.py:856 received :19199440.48:
2018-04-30 07:35:13.119Z root             10 feeControl.py:798 sending command :'~rt,ccd1\n':
2018-04-30 07:35:13.123Z root             10 feeControl.py:856 received :~rt,ccd1:
2018-04-30 07:35:13.790Z root             10 feeControl.py:856 received :18203293.12:
2018-04-30 07:35:13.791Z cmds             20 CommandLink.py:121 > 1 21590 i ccdTemps=270.63,19199440.48,18203293.12
2018-04-30 07:35:13.791Z cmds             20 CommandLink.py:121 > 1 21590 :

I feel that we have seen this before and that it was fixed (it must have been). But cannot find ticket/email.



 Comments   
Comment by cloomis [ 01/May/18 ]

We saw this before and did not fix it because 145K is 30K below operational temperatures, and there are other sensors near the CCD without this problem.

Comment by shope [ 01/May/18 ]

Measurements were below the physical measurement limits imposed by the hardware. All measurements are returned in Kelvin which means they cannot be negative. Firmware was modified to return -1.0 for measurements below the normal operating range.

Comment by cloomis [ 13/Jul/18 ]

See comments: This will not be an issue during operation, we do have other – not clipped – temperature sensors on the FPA, and the firmware now reports a sensibly invalid reading.

Generated at Sat Feb 10 16:24:00 JST 2024 using Jira 8.3.4#803005-sha1:1f96e09b3c60279a408a2ae47be3c745f571388b.